The sum of two consecutive numbers is -65. Find the value of the smaller number.
Found 2 solutions by math_helper, greenestamps:
Answer by math_helper(2461)


S + (S+1) = -65
2S + 1 = -65
2S = -66
S = -66/2
S = -33 --> S+1 = -32
The numbers -33 and -32 so the smaller is -33

Answer by greenestamps(13200)



Informally....

(-65)/2 = -32.5

so the two consecutive integers are -33 and -32.